1. Valles de Benasque y Eriste

ERISTE TO ANCILES THROUGH THE CONQUES TRACK 40 minutes

The departure point is the town of Eriste. Take the main road to Benasque but cross immediately to the right (near the lower end of the dam) the bridge over the Ésera river (to the left you go straight to the town of Anciles). You will find three forks, always take the tracks to the left till you get to the houses at the Conques. On their left you will find the Gardeta Fountain. The views of the valley are really beautiful from this spot.

Take the track next to the fountain to the left and you will get to Anciles. After a crossroad turn right, and after about 7 minutes (next to a house) leave the main track and take a narrower one on the right of the houses. After a while you will get to a wider track, which will take you to Anciles if you turn right and to Eriste if you turn left.